PR 25-5A Alternative capital investments Obj. 3, 4 projects, office expansion and The investment committee of Sentry Insurance Co. is evaluating two upgrade to computer servers. The projects have different useful lives, but each requires an invest- ment of $490,000. The estimated net cash flows from each project are as follows: 105 Net Cash Flows MPLATE Office Expansion Servers Year $125,000 $165,000 1 125,000 165,000 165,000 165,000 2 125,000 125,000 4 125,000 5 125,000 6 (Continued)
The committee has selected a rate of 12 % for purposes of net present value analysis. It also estimates that the residual value at the end of each project's useful life is $0, but at the end of the fourth year, the office expansion's residual value would be $180,000. Instructions 1. For each project, compute the net present value. Use the present value of an annuity of $1 table appearing in this chapter (Exhibit 5). (Ignore the unequal lives of the projects.) 2. For each project, compute the net present value, assuming that the office expansion is adjusted to a four-year life for purposes of analysis. Use the present value of $1 table appearing in this chapter (Exhibit 2). Prepare a report to the investment committee, providing your advice on the relative 3. merits of the two projects.
Answer #1
1) OFFICE EXPANSION:
NPV = -490000+125000*4.11141 = $ 23,926.25
SERVERS:
NPV = -490000+165000*3.03735 = $ 11,162.75
2) OFFICE EXPANSION:
NPV = -490000+125000*3.03735+180000*0.63552 = $    4,062.35
SERVERS:
NPV = -490000+165000*3.03735 = $ 11,162.75
3) Evaluated on a 4 year life basis, the project SERVERS has higher NPV
and is hence preferable.